OpenStudy (anonymous):

how do u multiply 3/10 * 90

OpenStudy (curry):

3 * 90 / 10

OpenStudy (anonymous):

3/10 * 90/1 = 270/10

OpenStudy (curry):

270 / 10 = 27

OpenStudy (anonymous):

= 27

OpenStudy (curry):

dont forget medal